OpenStudy (anonymous):

how to integrate e^2x ?

OpenStudy (anonymous):

\[\int\limits_{}^{}e^xdx = e^x + c\]

OpenStudy (anonymous):

\[\int\limits_{}^{}e^{2x}.dx = e^{2x}/2 + c\]

OpenStudy (anonymous):

yes, but how?

OpenStudy (anonymous):

explain.

OpenStudy (sriram):

differentiate rhs and see u will get the lhs

OpenStudy (anonymous):

rhs?? lhs??

OpenStudy (unklerhaukus):

you could substitute to be more explicit \[\alpha x=u\qquad\qquad\alpha \text d x=\text du\qquad\qquad \text dx=\frac{\text du}{\alpha}\] \[\int e^{\alpha x}\text dx=\int e^{u}\frac{\text du}{\alpha}=\frac 1\alpha \int e^u\text du=\frac 1 \alpha e^u+c=\frac{e^{\alpha x}}{\alpha}+c\]

OpenStudy (sriram):

UnkleRhaukus is right btw rhs and lhs are right hand side & left hand side